OpenStudy (anonymous):

There are red and blue marbles in a jar. One-third of the marbles are red. If another red marble is added to the jar, then 3/8 of the marbles will be red. How many marbles are blue?

hero (hero):

I guess 5/8 wasn't the correct answer

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

It needs to be a whole #

hero (hero):

Oh, I see

OpenStudy (anonymous):

x = Total number of marbles. y = Number of red initially. z = Number of blue initially. \[x/3 = y\] \[3x/8 = y + 1\] (Substitute)\[9y/8 = y + 1\]\[y = 8\]\[x = 24\]\[x = y + z\] \[z = 16\]

OpenStudy (anonymous):

I think thats right

OpenStudy (anonymous):

It's right lol :D

OpenStudy (anonymous):

thank u so much :D